In an independent probability, the probability of success is 0.43. There are 9 trials.

What is the probability of 1 success, 3 failures, 2 successes, and 3 failures in 9 trials, in that exact order?

(PLEASE WRITE ANSWER OUT)

Group of answer choices

0.0025

0.0030

0.0032

0.0022

0.0027
Answer #1

Solution:

Probability of success = p = 0.43

Probability of failures = q = 1 - 0.43 = 0.57

Number of trials = n = 9

Trials are independent.

P[ 1 success, 3 failures, 2 successes, and 3 failures ]

= P[1 success]* P[3 failures] * P[2 successes ] * P[3 failures]

= 0.431 * 0.573 * 0.432 * 0.573

= 0.43 * 0.185193 * 0.1849 * 0.185193

= 0.00272680763

= 0.0027

Answer : 0.0027